OpenStudy (anonymous):

Tricky one... Suppose that x* means 5x2 - x, then what does (y+3)* mean?

OpenStudy (anonymous):

it means replace x in the first expression by \[y+3\]

OpenStudy (anonymous):

i.e. you get \[(y+3)*=5(y+3)^2-(y+3)\] then you can clean it up with some algebra

OpenStudy (anonymous):

Oh wow! So, (y+3)* = 5y^2+30y+45? No further?

OpenStudy (anonymous):

But -(y+3) I don't know how to subtract that from it...

OpenStudy (anonymous):

Would (y+) and -(y+3) cross out?

OpenStudy (anonymous):

Please help!! I think I have it.. (y+3)* = 5y+15

Directrix (directrix):

(y+3)* = 5y^2+30y+45 - (y + 3) (y+3)* = 5y^2+30y+45 -1 (y + 3) (y+3)* = 5y^2+30y+45 -y -3 (y+3)* = 5y^2 +29y - 42